# PATS: Proficiency-Aware Temporal Sampling for Multi-View Sports Skill Assessment

### v3

@@ II-B Action Quality Assessment
Multi-modal approaches integrate skeleton features [13, 14] and audio information [15, 16] for enhanced understanding. Multi-view methods show particular promise: SkillFormer [1] processes synchronized egocentric and exocentric videos through a TimeSformer backbone with LoRA fine-tuning, using CrossViewFusion with multi-head cross-attention for view-specific feature fusion. However, it relies on uniform sampling, which fails to preserve complete temporal dynamics. EgoPulseFormer [17] leverages physiological signals from eye-tracking cameras. Datasets such as EgoExoLearn [18] and EgoExo4D [19] provide multi-perspective benchmarks.
@@ III-A Proficiency-Aware Temporal Sampling (PATS)
We introduce Proficiency-Aware Temporal Sampling (PATS), a novel temporal sampling strategy designed to preserve the sequential nature of skilled performances for accurate proficiency assessment. Unlike traditional uniform sampling methods that randomly select frames across a video, PATS extracts continuous temporal segments that maintain the natural flow of athletic movements and performance dynamics (1).
@@ III-B Integration with SkillFormer
We integrate PATS with SkillFormer [1] as it represents the current state-of-the-art for proficiency estimation on EgoExo4D, with both approaches targeting sports domains where temporal continuity distinguishes expert from novice performance.
PATS enhances SkillFormer’s multi-view architecture by preserving temporal continuity essential for cross-view fusion. SkillFormer’s CrossViewFusion module requires meaningful temporal relationships to capture performance nuances across camera perspectives. By providing continuous temporal segments rather than sparse frames, PATS enables more effective attention computations for multi-view movement analysis.
PATS seamlessly replaces SkillFormer’s original sampling while maintaining full architectural compatibility, enhancing proficiency assessment without requiring modifications.
@@ IV-B Implementation Details
To ensure fair comparison, we maintain identical training configurations to the original SkillFormer implementation [1]. All models are initialized from a TimeSformer backbone [20], pre-trained on Kinetics-600 [21], and fine-tuned for 4 epochs using AdamW optimizer with weight decay of 0.01 and LoRA adaptation. We adopt SkillFormer’s hyperparameters for each experimental setup.
@@ VI Limitations and Future Work
Future research should focus on automated configuration selection mechanisms, enhanced temporal modeling for rhythmic activities, and multi-modal integration incorporating audio and haptic feedback. Meta-learning approaches for rapid domain adaptation could further improve generalization across diverse skill assessment domains.

### v4

@@ II-B Action Quality Assessment
Multi-modal approaches integrate skeleton features [13, 14] and audio information [15, 16] for enhanced understanding. Multi-view methods show particular promise: SkillFormer [1] processes synchronized egocentric and exocentric videos through a TimeSformer [17] backbone with LoRA [18] fine-tuning, using CrossViewFusion with multi-head cross-attention for view-specific feature fusion. However, it relies on uniform sampling, which fails to preserve complete temporal dynamics. EgoPulseFormer [19] leverages physiological signals from eye-tracking cameras. Datasets such as EgoExoLearn [20] and EgoExo4D [21] provide multi-perspective benchmarks.
@@ III-A Proficiency-Aware Temporal Sampling (PATS)
We introduce Proficiency-Aware Temporal Sampling (PATS), a novel temporal sampling strategy designed to preserve the sequential nature of skilled performances for accurate proficiency assessment. Unlike traditional uniform sampling methods that randomly select frames across a video, PATS extracts continuous temporal segments that maintain the natural flow of athletic movements and performance dynamics, as demonstrated in Figure 1.
@@ III-B Integration with SkillFormer
We integrate PATS with SkillFormer [1] as it represents the current state-of-the-art for proficiency estimation on EgoExo4D [21], with both approaches targeting sports domains where temporal continuity distinguishes expert from novice performance.
PATS enhances SkillFormer’s multi-view architecture by preserving temporal continuity essential for cross-view fusion. The CrossViewFusion module relies on meaningful temporal relationships to capture performance nuances across camera perspectives. By providing continuous temporal segments rather than sparse frames, PATS enables more effective attention computations for multi-view movement analysis.
PATS integrates seamlessly into SkillFormer, replacing the original uniform sampling strategy while maintaining full architectural compatibility and enhancing proficiency assessment without structural modifications.
@@ IV-B Implementation Details
To ensure fair comparison, we maintain identical training configurations to the original SkillFormer implementation [1]. All models are initialized from a TimeSformer [17] backbone, pre-trained on Kinetics-600 [22], and fine-tuned for 4 epochs using AdamW optimizer with weight decay of 0.01 and LoRA adaptation. We adopt SkillFormer’s hyperparameters for each experimental setup.
@@ VI Limitations and Future Work
Future research should focus on automated configuration selection mechanisms, enhanced temporal modeling for rhythmic activities, and multi-modal integration incorporating audio and haptic feedback.